# Cinematic Gaussians: Real-Time HDR Radiance Fields with Depth of Field

### v2

@@ 1 Introduction
Radiance fields [MST21] remain the most prevalent representation for novel-view synthesis applications due to their outstanding ability to produce high-quality renderings of scenes from multi-view photos. 3D Gaussian representations [KKLD23] have further advanced the field and now routinely deliver real-time rendering performance. Most works in this space are concerned with accurate, all-in-focus reconstructions of well-lit scenes. However, for radiance fields to become versatile tools for storytelling and artistic exploration, i.e., to become cinematic, they must simultaneously handle high-contrast illumination and depth of field. To achieve this, we propose a method for reconstructing high-dynamic-range (HDR) radiance fields with real-time re-exposure and re-focusing capabilities based on multi-view low-dynamic-range (LDR) images with varying exposures, apertures, and focal distances.
@@
Considering the significance of the task, naturally, individual aspects of our problem setting have already been investigated in the literature. On the one hand, the limited dynamic range of typical cameras often leads to overexposed and underexposed regions, causing saturation (clamping to white) and noise, respectively [RHD10]. These limitations directly carry over to radiance fields based on such photos and significantly restrict post-editing capabilities. Addressing this issue, several works have considered reconstructing HDR radiance fields from multi-view exposure brackets [HZF22, JSYJYBO22, RFS22] or RAW data [MHMB22]. On the other hand, defocus blur in input images has been regarded as a degradation to overcome for all-in-focus reconstruction [WLP22, MLL22, LLSL23, LLS24], with synthetic depth of field as an occasional afterthought. In contrast to all these works, we embrace both HDR and depth of field as expressive tools and set out to develop a radiance field reconstruction approach that enables real-time rendering and high-quality post-editing of exposure, aperture, and focal distance.
@@
Input to our method is a set of LDR images that can be captured with a consumer-grade camera. Different from typical radiance field reconstruction settings, the input images vary not only in extrinsic camera parameters, but also in exposure time, aperture size, and focal distance. This diverse input sample distribution allows our method to build an all-in-focus HDR 3D Gaussian representation that is optimized for a subsequent application of lens blur. Depth of field is modeled as a highly efficient convolution of projected scene Gaussians with a Gaussian aperture function [KZB03], based on a thin-lens model. The convolved Gaussians are then rasterized and projected to LDR using a dedicated tonemapping module. Once trained, our model enables novel-view synthesis with arbitrary combinations of camera and lens parameters, facilitating flexible post-editing in terms of exposure and depth of field in real time. We show that our approach can synthesize high-quality and editable imagery, including appealing bokeh effects that require the joint treatment of HDR and depth of field. We also demonstrate that our system is on par with and often outperforms the state of the art in the individual disciplines of HDR and all-in-focus reconstruction. This evaluation is based on multiple new datasets we introduce, including both synthetic and real data, which we will make publicly available on publication. Please refer to the video presentation CinematicGS. For an enhanced visual experience, it is recommended to download and view it. Our contributions can be summarized as follows:
@@ 2.3 Depth-of-Field and High Dynamic Range Radiance Fields
Significant progress has been made in reconstructing radiance fields from 2D images, with NeRF [MST21] and 3D Gaussian Splatting [KKLD23] techniques being particularly noteworthy. Neural Radiance Fields (NeRF) [MST21] use a multi-layer perceptron (MLP) with positional encoding to map 3D locations and viewing directions to color and density values. Parameter optimization is achieved via gradient descent on the photometric error. Subsequent work on NeRF [HSM21, VHM22, BMT21, BMV22, BMV23, MESK22] focused on further improving its quality and initially slow rendering speed of this implicit representation. In contrast, 3D Gaussian Splatting (3DGS) [KKLD23] uses explicit 3D Gaussian primitives as a flexible and expressive scene representation. The optimization starts from point clouds and optimizes per-Gaussian properties. While effective, the recency of 3DGS entails minor flaws in its method. Mip-Splatting [YCH23] and multi-scale 3D Gaussian splatting [YLCL23] both address aliasing artifacts present in the original approach. Mip-Splatting uses 3D size constraints and 2D Mip filters to enhance image quality for zoomed-in or zoomed-out views. Multi-scale 3D Gaussian splatting maintains Gaussians at different scales, using more small Gaussians for high resolution and larger Gaussians for low resolution.
@@
In most radiance field research, it is assumed that input images are consistently in focus and well-exposed. While many methods can tolerate small deviations from those assumptions, few recent methods more actively model variations in focal length and exposure time. DoF-NeRF [WLP22] extends NeRF to handle shallow depth-of-field inputs and simulate DoF effects through a differentiable circle of confusion representation. Deblur-NeRF [MLL22] improves NeRF by addressing image blurriness from defocus or motion with a Deformable Sparse Kernel (DSK) module, enabling sharp 3D reconstructions. DP-NeRF [LLSL23] tackles geometric and appearance consistency issues in blurred images using physical priors and adaptive weighting, significantly enhancing 3D reconstruction quality in the presence of motion and defocus blur. Concurrent work [LLS24, DPRBK24] proposes to compensate for defocus blur in captures with Gaussian splats by inferring alternative covariances for out-of-focus Gaussians. HDR-NeRF [HZF22] recovers HDR radiance fields from LDR views with different exposures, enabling the generation of novel HDR and LDR views by modeling the physical imaging process integrated with NeRF. In freshly appearing concurrent work [CXL24], a similar approach has been proposed for 3DGS that greatly reduces training and inference time. RAWNeRF [MHMB22] utilizes a modified NeRF technique that trains directly on linear raw images to preserve the scene’s full dynamic range, enabling high-quality HDR novel view synthesis. These methods can handle simple DoF and HDR factors independently. However, they cannot account for both simultaneously; this causes issues in casual captures, e.g., scenes that contain details in the near foreground and far background, as well as strong luminance variation between dark shadows and bright lights.
@@ 3 Method
Our approach takes as input a set of LDR multi-view images $\{I_{k}\}_{k=1}^{N}$ of a scene, captured with randomly selected exposure times $t$ , aperture sizes $a$ , and focal distances ${d_{f}}$ (Fig. Cinematic Gaussians:  Real-Time HDR Radiance Fields with Depth of Field, left). Since exposure and lens settings are typically stored in EXIF headers, we assume this information is available for each image. Using this input, our method reconstructs an HDR radiance field that allows for post-editing of exposure and depth of field.
@@ 3.2 Camera Calibration
We require posed input images to build our model. The standard approach for this is Structure from Motion (SfM) [SF16], which we run on our input data. Unfortunately, such photometric calibration leaves the scene’s global scale ambiguous. As various lens parameters are provided in absolute millimeters, we need to scale the scene accordingly. To achieve this, we leverage the defocused images and their respective lens configurations. Specifically, we run an image blur detector [AGK17] on all input images to obtain per-pixel sharpness estimates. By thresholding these sharpness maps, we identify the image space locations of in-focus scene elements. The known focal distance ${d_{f}}$ per input image provides ground-truth depth values for these pixels. Next, we train a vanilla 3D Gaussian Splatting model on the input images after rescaling pixel intensities to a common exposure. Using this model, we render a depth map for each input view. As expected, the renderings from this model are of low quality given our inputs, but we find the overall scene configuration precise enough for our calibration. Using the depth values of the in-focus pixels, we perform a least-squares fit of the scene scale to match the rendered depth values with the ground truth.
After calibration, we have the full set of camera and lens parameters per input view at our disposal.
@@ 3.3 HDR Radiance Field with Depth of Field
where $f$ is the focal length, $a$ is the aperture size, and ${d_{f}}$ is the focal distance. We seek to enlarge each projected scene Gaussian to reflect the circle of confusion [KZB03], where the mean $\boldsymbol{\mu}$ determines depth 0pt. We approximate the circle of confusion using an aperture Gaussian with covariance $\hat{\Sigma}^{a}=\text{diag}\left(3r(0pt),3r(0pt)\right)\in\mathds{R}^{2\times 2}$ to allow for efficient convolutions with the scene Gaussians. This convolution can be carried out in closed form [Bro03, CHK22], resulting in an update of Eq. 2:
@@
simulating the per-pixel non-linear response of a real-world camera [AT11]. Here, the final exposure is determined using exposure time $t$ and aperture size $a$ , before the intensities are clipped to the target domain. Finally, $\gamma$ performs a gamma correction from linear radiance to pixel intensities.
Eq. 7 produces images that the user can flexibly edit across multiple dimensions: In addition to free viewpoint selection, the user can adjust the aperture size $a$ , exposure time $t$ , and focal distance ${d_{f}}$ in real-time, providing a wide range of artistic control.
@@ 3.4 Training
We set $\lambda_{\text{exp}}=0.25$ and $\lambda_{\text{foc}}=1$ in all our experiments and train our models using the Adam [KB15] optimizer with default parameters.
@@ 4.1.1 Dataset
Synthetic dataset. Our synthetic dataset is composed of four diverse indoor scenes rendered with Cycles (path tracer) in Blender 4.1. The scenes encompass challenging scenarios with high dynamic range of illumination (bright light sources and dark regions illuminated by bounced light) and spacious layouts allowing for multiple focal planes (examples shown in Figures 4 and 6). All scenes have been rendered with high number of samples (ranging from 4096 to 16384) to prevent undesired distortions in the defocused regions caused by denoising. For each scene we generated 960 HDR images (8 $\times$ 8 views sampled at a regular grid wrapped to the sphere surface, 5 uniformly sampled focus distances and 3 different apertures). For each view we generated all-in-focus ground truth image. This results in 1024 images per scene. For evaluation purposes we generated another set of 7 $\times$ 7 views located between the training views. All images are accompanied by a file containing parameters used for rendering (camera transformation matrix, focal length, focus distance, aperture, etc.). In addition to the HDR content we have modified the scenes for the AiF reconstruction comparison, producing pleasant LDR images with reduced dynamic range without clipped regions (examples shown in Figure 3). The camera locations, apertures, and focus distances are sampled in a same manner as in the case of the HDR subset.
Captured dataset. Our captured dataset is composed of four real scenes. Some examples are shown in Figs. Cinematic Gaussians:  Real-Time HDR Radiance Fields with Depth of Field, 5, and 6. These scenes were captured using Canon EOS RP and Canon EOS Rebel T6i cameras, outfitted with RF 85mm F2.0, and EF-S 18-55mm F3.5-5.6 lenses. Each scene was captured by moving the camera to follow an approximately 9$\times$9 grid pattern, resulting in 81 images per scene. At each camera position in this grid, we varied the exposure time, aperture, and focus distance to ensure a diverse set of configurations.
@@ 4.1.2 Implementation
We use the 3DGS framework [KKLD23] as our baseline, integrating a thin-lens model and a tone mapping module as described in Sec. 3.3. The implementation is done in PyTorch and CUDA. We will release all source code on publication.
Additionally, we incorporate some practical design choices to tackle our task more effectively. Given that our rendering dataset starts from random noise without an SfM point cloud initialization, we employ a coarse-to-fine optimization strategy. During the initial 7,000 iterations, we exclude tone mapping and depth of field simulation, training on an exposure-aligned dataset to generate an initial point cloud. Following this, we refine the optimization using the coarse result as a starting point. The iteration process consists of 40,000 steps, taking approximately 21 minutes on an RTX A40 GPU, using a training dataset with a resolution of 1200 × 675.
@@ 4.2.1 All-in-Focus Reconstruction
For the all-in-focus reconstruction task, we compare our results against the concurrent work Deblur-Splatting [LLS24]. Additionally, we include in the supplementary material comparisons for Deblur NeRF [MLL22], where we observed noticeable deformations along the boundaries, causing straight lines to appear curved. We use the four scenes from our synthetic dataset for this evaluation. In particular, for each of the scenes, we use 64 (8 $\times$ 8) viewpoints for training, each exhibiting varying degrees of defocus blur, and 49 (7 $\times$ 7) novel viewpoints not seen during training for the evaluation. Table 1 presents the numerical results for PSNR, SSIM, and LPIPS, commonly used for evaluating image quality. Additionally, Fig. 3 shows illustrative comparisons to the ground truth. Although our method is not specifically designed for deblurring, it achieves results comparable to state-of-the-art method. Regarding rendering time, both methods are similarly efficient, achieving approximately 110 frames per second (FPS) on an RTX A40 GPU.
Table 1: Comparison of all-in-focus reconstruction performance metrics (PSNR, SSIM, and LPIPS) for our method and Deblur-Splatting using our synthetic dataset.

### v3

@@ Abstract
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010382.10010236</concept_id> <concept_desc>Computing methodologies Computational photography</concept_desc> <concept_significance>300</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293.10010294</concept_id> <concept_desc>Computing methodologies Neural networks</concept_desc> <concept_significance>300</concept_significance> </concept> <concept>
<concept> <concept_id>10010147.10010371.10010382.10010385</concept_id> <concept_desc>Computing methodologies Image-based rendering</concept_desc> <concept_significance>300</concept_significance> </concept> </ccs2012>
</ccs2012>
@@ 1 Introduction
Radiance fields [MST21] remain the most prevalent representation for novel-view synthesis applications due to their outstanding ability to produce high-quality renderings of scenes from multi-view photos. 3D Gaussian representations [KKLD23] have further advanced the field and now routinely deliver real-time rendering performance. Most works in this space are concerned with accurate, all-in-focus reconstructions of well-lit scenes. However, for radiance fields to become versatile tools for storytelling and artistic exploration, i.e., to become cinematic, they must simultaneously handle high-contrast illumination and depth of field. To achieve this, we propose a method for reconstructing high-dynamic-range (HDR) radiance fields with real-time re-exposure and re-focusing capabilities based on multi-view low-dynamic-range (LDR) images with varying exposures, apertures, and focus distances.
@@
Considering the significance of the task, naturally, individual aspects of our problem setting have already been investigated in the literature. On the one hand, the limited dynamic range of typical cameras often leads to overexposed and underexposed regions, causing saturation (clamping to white) and noise, respectively [RHD10]. These limitations directly carry over to radiance fields based on such photos and significantly restrict post-editing capabilities. Addressing this issue, several works have considered reconstructing HDR radiance fields from multi-view exposure brackets [HZF22, JSYJYBO22, RFS22] or RAW data [MHMB22]. On the other hand, defocus blur in input images has been regarded as a degradation to overcome for all-in-focus reconstruction [WLP22, MLL22, LLSL23, LLS24], with synthetic depth of field as an occasional afterthought. In contrast to all these works, we embrace both HDR and depth of field as expressive tools and set out to develop a radiance field reconstruction approach that enables real-time rendering and high-quality post-editing of exposure, aperture, and focus distance.
@@
The input to our method is a set of LDR images that can be captured with a consumer-grade camera. Different from typical radiance field reconstruction settings, the input images vary not only in extrinsic camera parameters, but also in exposure time, aperture size, and focus distance. This diverse input sample distribution allows our method to build an all-in-focus HDR 3D Gaussian representation that is optimized for a subsequent application of lens blur. Depth of field is modeled as a highly efficient convolution of projected scene Gaussians with a Gaussian aperture function [KZB03], based on a thin-lens model. The convolved Gaussians are then rasterized and projected to LDR using a dedicated tonemapping module. Once trained, our model enables novel-view synthesis with arbitrary combinations of camera and lens parameters, facilitating flexible post-editing in terms of exposure and depth of field in real time. We show that our approach can synthesize high-quality and editable imagery, including appealing bokeh effects that require the joint treatment of HDR and depth of field. We also demonstrate that our system is on par with and often outperforms the state of the art in the individual disciplines of HDR and all-in-focus reconstruction. This evaluation is based on multiple new datasets we introduce, including both synthetic and real data, all available on our webpage https://cinegs.mpi-inf.mpg.de/.
Our contributions can be summarized as follows:
@@ 2.3 Depth-of-Field and High Dynamic Range Radiance Fields
Significant progress has been made in reconstructing radiance fields from 2D images, with NeRF [MST21] and 3D Gaussian Splatting [KKLD23] techniques being particularly noteworthy. Neural Radiance Fields (NeRF) [MST21] use a multi-layer perceptron (MLP) with positional encoding to map 3D locations and viewing directions to color and density values. Parameter optimization is achieved via gradient descent on the photometric error. Subsequent work on NeRF [HSM21, VHM22, BMT21, BMV22, BMV23, MESK22] focused on further improving its quality and initially slow rendering speed of this implicit representation. In contrast, 3D Gaussian Splatting (3DGS) [KKLD23] uses explicit 3D Gaussian primitives as a flexible and expressive scene representation. The optimization starts from point clouds and optimizes per-Gaussian properties. While effective, the recency of 3DGS entails minor flaws in its method. Mip-Splatting [YCH24] and multi-scale 3D Gaussian splatting [YLCL24] both address aliasing artifacts present in the original approach. Mip-Splatting uses 3D size constraints and 2D Mip filters to enhance image quality for zoomed-in or zoomed-out views. Multi-scale 3D Gaussian splatting maintains Gaussians at different scales, using more small Gaussians for high resolution and larger Gaussians for low resolution.
@@
In most radiance field research, it is assumed that input images are consistently in focus and well-exposed. While many methods can tolerate small deviations from those assumptions, few recent methods more actively model variations in focal length and exposure time. DoF-NeRF [WLP22] extends NeRF to handle shallow depth-of-field inputs and simulate DoF effects through a differentiable circle of confusion representation. Deblur-NeRF [MLL22] improves NeRF by addressing image blurriness from defocus or motion with a Deformable Sparse Kernel (DSK) module, enabling sharp 3D reconstructions. DP-NeRF [LLSL23] tackles geometric and appearance consistency issues in blurred images using physical priors and adaptive weighting, significantly enhancing 3D reconstruction quality in the presence of motion and defocus blur. Concurrent work [LLS24, DPRBK24] proposes to compensate for defocus blur in captures with Gaussian splats by inferring alternative covariances for out-of-focus Gaussians. HDR-NeRF [HZF22] recovers HDR radiance fields from LDR views with different exposures, enabling the generation of novel HDR and LDR views by modeling the physical imaging process integrated with NeRF. In freshly appearing concurrent work [CXL24], a similar approach has been proposed for 3DGS that greatly reduces training and inference time. RawNeRF [SGHS98, MHMB22] utilizes a modified NeRF technique that trains directly on linear raw images to preserve the scene’s full dynamic range, enabling high-quality HDR novel view synthesis. It should be noted that RawNeRF uses raw format inputs without defocus blur and requires the multiplane image-based blurring algorithm [KS07] for refocusing. In contrast, our approach uses defocused 8-bit LDR inputs, jointly models depth of field and tone mapping processes, and inherently supports all post-editing tasks. The methods discussed above can handle simple DoF and HDR factors independently. However, they cannot account for both simultaneously; this causes issues in casual captures, e.g., scenes that contain details in the near foreground and far background, as well as strong luminance variation between dark shadows and bright lights.
@@ 3 Method
Our approach takes as input a set of LDR multi-view images $\{I_{k}\}_{k=1}^{N}$ of a scene, captured with randomly selected exposure times $t$ , aperture sizes $a$ , and focus distances ${d_{f}}$ (Fig. Cinematic Gaussians:  Real-Time HDR Radiance Fields with Depth of Field, left). Since exposure and lens settings are typically stored in EXIF headers, we assume this information is available for each image. Using this input, our method reconstructs an HDR radiance field that allows for post-editing of exposure and depth of field.
@@ 3.2 Camera Calibration
We require posed input images to build our model. The standard approach for this is Structure from Motion (SfM) [SF16], which we run on our input data. Unfortunately, such photometric calibration leaves the scene’s global scale ambiguous. As various lens parameters are provided in absolute millimeters, we need to scale the scene accordingly. To achieve this, we leverage the defocused images and their respective lens configurations. Specifically, we run an image blur detector [AGK17] on all input images to obtain per-pixel sharpness estimates. By thresholding these sharpness maps, we identify the image space locations of in-focus scene elements. The known focus distance ${d_{f}}$ per input image provides ground-truth depth values for these pixels. Next, we train a vanilla 3D Gaussian Splatting model on the input images after rescaling pixel intensities to a common exposure. Using this model, we render a depth map for each input view. As expected, the renderings from this model are of low quality given our inputs, but we find the overall scene configuration precise enough for our calibration. Using the depth values of the in-focus pixels, we perform a least-squares fit of the scene scale to match the rendered depth values with the ground truth.
After calibration, we have the full set of camera and lens parameters per input view at our disposal. Please note that we follow the methodology established in 3DGS  [KKLD23]. The initialization process differs for synthetic and real datasets: for real scenes, we use SfM for camera and scale calibration, and initialize the scene Gaussians using SfM points. For synthetic scenes with predefined camera parameters, SfM is unnecessary, eliminating scale ambiguity and calibration. In this case, scene Gaussians are initialized with random noise (see Sec.  4.1.2).
@@ 3.3 HDR Radiance Field with Depth of Field
where $f$ is the focal length, $a$ is the F-number, representing the aperture size and ${d_{f}}$ is the focus distance. We seek to enlarge each projected scene Gaussian to reflect the circle of confusion [KZB03], where the mean $\boldsymbol{\mu}$ determines depth 0pt. We approximate the circle of confusion using an aperture Gaussian with covariance $\hat{\Sigma}^{a}=\text{diag}\left(3r(0pt),3r(0pt)\right)\in\mathds{R}^{2\times 2}$ to allow for efficient convolutions with the scene Gaussians. This convolution can be carried out in closed form [Bro03, CHK22], resulting in an update of Eq. 2:
@@
simulating the per-pixel non-linear response of a real-world camera [AT11]. Here, the final exposure is determined using exposure time $t$ and aperture size $a$ , before the intensities are clipped to the target domain. Finally, $\gamma$ performs a gamma correction from linear radiance to pixel intensities. We use an explicit tone mapping operator as it minimizes computational overhead and allows convenient metadata incorporation.
Eq. 7 produces images that the user can flexibly edit across multiple dimensions: In addition to free viewpoint selection, the user can adjust the aperture size $a$ , exposure time $t$ , and focus distance ${d_{f}}$ in real-time, providing a wide range of artistic control.
@@ 3.4 Training
We set $\lambda_{\text{exp}}=0.25$ and $\lambda_{\text{foc}}=1$ in all our experiments and train our models using the Adam [KB15] optimizer with default parameters, also including all densification strategies from 3DGS [KKLD23].
@@ 4.1.1 Dataset
Synthetic dataset Our synthetic dataset is composed of four diverse indoor scenes rendered with Cycles (path tracer) in Blender 4.1. The scenes encompass challenging scenarios with high dynamic range of illumination (bright light sources and dark regions illuminated by bounced light) and spacious layouts allowing for multiple focal planes (examples shown in Figures 4 and 6). All scenes have been rendered with high number of samples (ranging from 4096 to 16384) to prevent undesired distortions in the defocused regions caused by denoising. For each scene we generated 960 HDR images (8 $\times$ 8 views sampled at a regular grid wrapped to the sphere surface, 5 uniformly sampled focus distances and 3 different apertures). For each view we generated all-in-focus ground truth image. This results in 1024 images per scene. For evaluation purposes we generated another set of 7 $\times$ 7 views located between the training views. All images are accompanied by a file containing parameters used for rendering (camera transformation matrix, focal length, focus distance, aperture, etc.). In addition to the HDR content we have modified the scenes for the AiF reconstruction comparison, producing pleasant LDR images with reduced dynamic range without clipped regions (examples shown in Figure 3). The camera locations, apertures, and focus distances are sampled in a same manner as in the case of the HDR subset.
Captured dataset Our captured dataset is composed of four real scenes. Some examples are shown in Figs. Cinematic Gaussians:  Real-Time HDR Radiance Fields with Depth of Field, 5, and 6. These scenes were captured using Canon EOS RP and Canon EOS Rebel T6i cameras, outfitted with RF 85mm F2.0, and EF-S 18-55mm F3.5-5.6 lenses. Each scene was captured by moving the camera to follow an approximately 9$\times$9 grid pattern, resulting in 81 images per scene. At each camera position in this grid, we varied the exposure time, aperture, and focus distance to ensure a diverse set of configurations.
@@ 4.1.2 Implementation
We use the 3DGS framework [KKLD23] as our baseline, integrating a thin-lens model and a tone mapping module as described in Sec. 3.3. The implementation is done in PyTorch and CUDA. All source code is available on our project webpage.
Additionally, we incorporate some practical design choices to tackle our task more effectively. Given that our synthetic dataset starts from random noise without an SfM point cloud initialization, we employ a coarse-to-fine optimization strategy. During the initial 7,000 iterations, we exclude tone mapping and depth of field simulation, training on an exposure-aligned dataset to generate an initial point cloud. Following this, we refine the optimization using the coarse result as a starting point. Our first stage training adapts well to a broad range of dynamic variations. However, due to sparse viewpoints, we observed instability when the scene’s dynamic range exceeds 18 stops. The iteration process consists of 40,000 steps, taking approximately 21 minutes on an RTX A40 GPU, using a training dataset with a resolution of 1200 × 675.
@@ 4.2.1 All-in-Focus Reconstruction
For the all-in-focus reconstruction task, we compare our results against the concurrent work Deblur-Splatting [LLS24]. Additionally, we include in the supplementary material comparisons for Deblur NeRF [MLL22], where we observed noticeable deformations along the boundaries, causing straight lines to appear curved. We use the four scenes from our synthetic dataset for this evaluation. In particular, for each of the scenes, we use 64 (8 $\times$ 8) viewpoints for training, each exhibiting varying degrees of defocus blur, and 49 (7 $\times$ 7) novel viewpoints not seen during training for the evaluation. Table 1 presents the numerical results for PSNR, SSIM [WBSS04], and LPIPS [ZIE18], commonly used for evaluating image quality. Additionally, Fig. 3 shows illustrative comparisons to the ground truth. Although our method is not specifically designed for deblurring, it achieves results comparable to state-of-the-art method. Regarding rendering time, both methods are similarly efficient, achieving approximately 110 frames per second (FPS) on an RTX A40 GPU. We primarily evaluate the synthetic dataset due to the challenge of obtaining ground truth for real datasets. To validate our method, we run Deblur-Splatting on our real dataset using non-reference metrics. Since Deblur-Splatting cannot generate HDR, we tone-map our HDR images and evaluate in the LDR domain using NIQE [MSB12] and BRISQUE [MMB12]. The results in Table 2 demonstrate our method’s superior performance.
Table 1: Comparison of all-in-focus reconstruction performance metrics for our method and Deblur-Splatting using our synthetic dataset.